is a velocity selector that can be used to measure the speed of a charged particle. A beam of particles is directed along the axis of the instrument. A parallel plate capacitor sets up an electric field E which is oriented perpendicular to a uniform magnetic field B. If the plates are separated by 8 mm and the value of the magnetic field is 0.2 T, what voltage between the plates will allow particles of speed to pass straight through without deflection?


Tube (cutawayx x » x * view B (into paper) x x X3 x x xX x % x x Capacitor * * * plates x XX
Answer #1

Given that particle passes through without deflection, which means

Electrostatic force = Magnetic force

Fe = Fm

q*E = q*v*B*sin theta

theta = 90 deg

v = speed of particle

E = electric field = V/d

B = 0.2 T

d = separation between plates = 8 mm = 0.008 m

So,

q*V/d = q*v*B*sin 90 deg

V = v*B*d

V = v*0.2*0.008

V = 1.6*10^-3*v

See that value of velocity of particle should have been given in the question, check the question again and let me know in the comment section.

Please Upvote.